# On a pinning model in correlated Gaussian random environments

### v1

@@ 3.2. Some discussion on uniform integrability
The Weinrib-Halperin prediction suggests that our pinning model should be disorder relevant under (A1), i.e., $H\in(\frac{1}{2},1)$ and $\alpha+H>1$. This is partially confirmed by our results in the following sense: Proposition 3.1 yields the weak convergence of each chaos under (A1); while for the weak convergence of the partition functions, an extra condition $\alpha>\frac{1}{2}$ is imposed in order to guarantee the $L^{2}$-convergence of the Wiener chaos expansion (see Step (1) in the proof of Theorem 1.1 and Remark 3.3). This implies that when $H\in(\frac{1}{2},1)$ and $\alpha>\frac{1}{2}$, the family $\{\tilde{Z}_{N}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N}}\}_{N\in\mathbb{N}}$ of the rescaled partition functions is bounded in $L^{2}$-norm and in particular, for the limiting distribution we have $\mathbb{E}[\tilde{Z}_{\hat{\beta},\hat{h}}]=1$. Inspired by the Weinrib-Halperin prediction, we conjecture that the family $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is uniformly integrable and converges weakly to an $L^{1}$-integrable limit under (A1).
In the rest of this section, we shall prove the uniform integrability of $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ and characterize its weak limit, under a stronger condition:
@@
$\sup_{N}\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}<\infty,$ (3.15)
and
$\sup_{N}\mathbb{E}\Big{[}\exp\Big{\{}h_{N}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}\Big{\}}\Big{]}<\infty.$ (3.16)
@@
We prove the result for the case $\alpha<1$ and the other cases can be proven in a similarly way. For (3.15), we have
$\quad\beta_{N}^{2}\mathbb{E}\Big{[}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big{]}$
$=\hat{\beta}^{2}\sum_{n,m=1}^{N}N^{2-2H}\gamma(n-m)\frac{L(N)^{2}}{N^{2{\alpha}-2}}{\mathbb{P}}_{\tau}(n\in\tau,m\in\tau)\frac{1}{N^{2}}$
@@
for $\alpha+2H>2$, where $\phi_{2}$ is defined in (1.12) and the inequality follows from (1.6), (1.7) and (3.12). For (3.16), we have that for some $\alpha^{\prime}\in(0,\alpha)$
$\begin{split}\mathbb{E}\Big{[}\exp\Big{\{}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{\}}\Big{]}&=\sum\limits_{k=0}^{\infty}\frac{1}{k!}(h_{N})^{k}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{k}}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{k}\in\tau)\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{\Gamma(k\alpha^{\prime})}<\infty,\end{split}$ (3.17)
where we have used (3.12) and Lemma A.5. ∎
@@
Recall that
$\tilde{Z}_{N}=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{\}}\exp\Big{\{}\diamond\Big{(}\beta_{N}\sum_{x\in\Omega_{N}}\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]},$
and we have
@@
×$\quad\times\exp\Big{\{}h_{N}\sum_{n=1}^{N}1_{\{n\in\tau^{\prime}\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big{\}}\bigg{]}.$
@@
Lemma 3.4 and Markov’s inequality together yields:
@@
Then, by Lemma A.7, we get
$\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|\leq\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c}),$
where we denote
@@
$\mathbb{E}_{\omega}\left|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})\right|$ $\leq\mathbb{E}_{\omega}\mathbb{E}_{\tau}\left[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\right]+{\mathbb{P}}((A_{l}^{N})^{c})$ (3.19)
$=\mathbb{E}_{\tau}\Big{[}\mathbf{1}_{(A_{l}^{N})^{c}}\exp\Big{(}h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}+{\mathbb{P}}((A_{l}^{N})^{c})$
$\leq{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}\mathbb{E}_{\tau}\Big{(}\Big{[}\exp\Big{(}2h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big{)}\Big{]}\Big{)}^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c})$
$\leq C{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c}),$
@@
A direct consequence of the uniform integrability is the following.
@@
In the proof of Proposition 3.5, we apply Cauchy-Schwartz inequality to decouple $\tau$ and $\tau^{\prime}$, which requires the stronger Assumption (A2). We believe this is a technical issue and (A1) should be sufficient for the uniform integrability.
When $\alpha<\frac{1}{2}$, though $\{\tilde{Z}_{N}\}_{N\in\mathbb{N}}$ is tight, we are not able to prove the weak convergence as in the proof of Theorem 1.1. This is because, when $\alpha<\frac{1}{2}$, $\tilde{Z}_{N}$ is not uniformly bounded in $L^{2}$ (see Remark 3.3) and thus $\mathbb{E}[\mathcal{Z}^{2}]=\infty$ for any weak limit $\mathcal{Z}$. Therefore, we cannot employ the theory of Wiener chaos, which applies only to $L^{2}$-integrable random variables, to obtain a weak limit.
@@ 4. The Stratonovich case
[…]
@@
Assume (A2). Then we have $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $j=0,1,...,\lfloor\frac{m}{2}\rfloor$. Moreover, we have
@@
$\|\mathrm{Tr}^{j}{\phi}_{m}$ ||($\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\phi_{m}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
×$\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
First, we split $[0,1]^{2m}$ into $(2m)!$ regions such that in each region, the numbers $\{t_{i},t_{i}^{\prime},s_{j},s_{j}^{\prime}:i=1,\dots,m-2j,j=1,\dots,2j\}$ are listed in an increasing order. To get an upper bound, we replace, for instance, each $|t_{i}-t_{i}^{\prime}|^{2H-2}$ by $|t_{i}^{*}-t_{i}^{\star}|^{2H-2}$ where $t^{*}_{i}:=\max\{t_{i},t_{i}^{\prime}\}$ and $t_{i}^{\star}$ is the largest number that is smaller than $t^{*}_{i}$ (it could be some $t_{j}$, $t^{\prime}_{j}$, $s_{j}$ or $s^{\prime}_{j}$), and we perform the same operation for all the other $|s_{2i}-s_{2i-1}|^{2H-2}$, $|s^{\prime}_{2i}-s^{\prime}_{2i-1}|^{2H-2}$ and the factors in $\phi_{m}$ (see Figures 1 and 2 for illustration). Then on each region, the above integral is smaller than an integral with $2m$ variables in Lemma A.5 with $m$ exponents equal to ${\alpha}-1$ and $m$ exponents equal to $32H+{\alpha}-$ 3. Noting that $12H+\alpha-3>-$ 1 under (A2), we get by Lemma A.5
@@
[…]
Figure 1. An example for $m=6$ and $j=2$ before modifications, the red segments represent the renewal kernels (with exponent $\alpha-1$) and the blue segments represent the kernels of the fractional Brownian motion (with exponent $22H-$ 2).
[…]
Figure 2. After modifications, the red segments represent the renewal kernels and the blue segments represent the kernels of the fractional Brownian motion.
When $k>0$, we have
$\quad\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}$
$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg{(}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
×$\quad\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}$
×$\qquad\times\bigg{(}\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\mathrm{d}\boldsymbol{x}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
×$\qquad\times\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\mathrm{d}\boldsymbol{x^{\prime}}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$\leq\int_{[0,1]^{2(m+k-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\prod_{p=1}^{k}|x_{p}-x_{p}^{\prime}|^{2H-2}\mathrm{d}x_{p}\mathrm{d}x_{p}^{\prime}$
×$\quad\times\bigg{(}\int_{[0,1]^{2j}}\phi_{m+k}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j},x_{1},\cdots,x_{k})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
×$\quad\times\int_{[0,1]^{2j}}\phi_{m+k}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime},x_{1}^{\prime},\cdots,x_{k}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg{)}$
$=\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$
@@
Assuming (A2), Lemma 4.1 yields ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$ (see (2.9)) and hence $\mathbb{I}_{m}(\psi_{m,k})$ is well defined due to Hu-Meyer’s formula (2.8). On the other hand, $\alpha+2H>2$ is the optimal condition for ${\vvvert\psi_{m,k}\vvvert_{m}<\infty}$. Indeed, for $k=0,m=2$ and $j=1$, we have
[…]
@@
Recalling
$Z_{\hat{\beta},\hat{h}}=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\mathbb{I}_{m}(\psi_{m,k}),$
we have
$\|Z_{\hat{\beta},\hat{h}}\|_{2}$ $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\|\mathbb{I}_{m}(\psi_{m,k})\|_{2}$
$\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}\hat{h}^{k}}{m!k!}\sqrt{2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}}$
@@
where the second inequality follows from Lemma 4.4 below and the last inequality is due to (4.2) and (4.3) below.
@@
$\mathbb{E}[\mathbb{I}_{m}(\phi_{m})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\phi}_{2m},$
and
$\mathbb{E}[\mathbb{I}_{m}(\psi_{m,k})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}.$
@@
For the first inequality, Hu-Meyer’s formula (2.8) implies
$\mathbb{E}[|\mathbb{I}_{m}(\phi_{m})|^{2}]$ $=\sum_{k=0}^{[\frac{m}{2}]}\frac{1}{(m-2k)!}\left(\frac{m!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.4)
$=\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}\binom{m}{2k}\binom{2k}{k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}\leq\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}2^{m+2k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$
$=2^{m}m!\sum_{k=0}^{[\frac{m}{2}]}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}},$

### v2

@@ 3.2. Some discussion on uniform integrability
[…]
Considering the above discussion together with Remarks 1.6 and 3.3, we conjecture that the family $\{\tilde{Z}_{N}\}_{N}$ converges weakly to some $L^{1}$-integrable random variable under (A1). In the rest of this section, we shall prove the uniform integrability of $\{\tilde{Z}_{N}\}_{N}$ and characterize its weak limit, under a stronger condition:
@@
$\sup_{N}~\beta_{N}^{2}\mathbb{E}\Big[\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big]<\infty\quad\text{and}\quad\sup_{N}~\mathbb{E}\Big[\exp\Big\{h_{N}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}\Big\}\Big]<\infty.$ (3.18)
@@
We prove the result for the case $\alpha<1$ and the other cases can be proven in a similar way. For the first bound in (3.18), we have
$\quad\beta_{N}^{2}\mathbb{E}\Big[\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau\}}\mathbf{1}_{\{m\in\tau\}}\Big]$ $=\hat{\beta}^{2}\sum_{n,m=1}^{N}N^{2-2H}\gamma(n-m)\frac{L(N)^{2}}{N^{2{\alpha}-2}}{\mathbb{P}}_{\tau}(n\in\tau,m\in\tau)\frac{1}{N^{2}}$
@@
for $\alpha+2H>2$, where $\phi_{2}$ is defined in (1.13) and the inequality follows from (1.6), (1.7) and (3.13). For the second bound in (3.18), we have that for some $\alpha^{\prime}\in(0,\alpha)$
$\begin{split}\mathbb{E}\Big[\exp\Big\{\sum_{n=1}^{N}h_{N}\mathbf{1}_{\{n\in\tau\}}\Big\}\Big]&=\sum\limits_{k=0}^{\infty}\frac{1}{k!}(h_{N})^{k}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{k}}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{k}\in\tau)\\ &=\sum\limits_{k=0}^{\infty}\frac{\hat{h}^{k}}{k!}\int_{[0,1]^{k}}g_{0,k}^{(N)}(t_{1},\cdots,t_{k})\prod\limits_{i=1}^{k}{\mathrm{d}}t_{i}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{\Gamma(k\alpha^{\prime})}<\infty.\end{split}$ (3.19)
where we have used (3.10), (3.13), (1.13) and Lemma A.5. ∎
@@
Recall (1.11) and we have
@@
$\quad\times\exp\Big\{h_{N}\sum_{n=1}^{N}1_{\{n\in\tau^{\prime}\}}+\frac{\beta_{N}^{2}}{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n\in\tau^{\prime}\}}\mathbf{1}_{\{m\in\tau^{\prime}\}}\Big\}\bigg].$
@@
Lemma 3.4 and Markov’s inequality together yield:
@@
Then, by Lemma A.7, we get $|\tilde{Z}_{N}-\tilde{Z}_{N}(A_{l}^{N})|\leq\mathbb{E}_{\tau}[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}]+{\mathbb{P}}((A_{l}^{N})^{c})$, where
@@
$\mathbb{E}_{\omega}\big|\tilde{Z}_{N}$- $\tilde{Z}_{N}(A_{l}^{N})\big|\leq\mathbb{E}\big[\mathscr{E}_{N}\mathbf{1}_{(A_{l}^{N})^{c}}\big]+{\mathbb{P}}((A_{l}^{N})^{c})=\mathbb{E}_{\tau}\Big[\mathbf{1}_{(A_{l}^{N})^{c}}\exp\Big(h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big)\Big]+{\mathbb{P}}((A_{l}^{N})^{c})$ (3.21)
$\leq{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}\mathbb{E}_{\tau}\Big(\Big[\exp\Big(2h_{N}\sum_{x\in\Omega_{N}}\delta_{x}\Big)\Big]\Big)^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c})\leq C{\mathbb{P}}((A_{l}^{N})^{c})^{\frac{1}{2}}+{\mathbb{P}}((A_{l}^{N})^{c}),$
@@
In the proof of Proposition 3.5, we applied the Cauchy-Schwarz inequality to decouple $\tau$ and $\tau^{\prime}$, which requires the stronger Assumption (A2). It seems that this is a technical issue and (A1) should be sufficient for the uniform integrability. A direct consequence of the uniform integrability is the following, which characterizes the possible weak limit of $\{\tilde{Z}_{N}\}_{N}$.
@@ 4. The Stratonovich case
$=\sum_{k=0}^{\infty}\sum_{p+q=k}\frac{\hat{\beta}^{p}\hat{h}^{q}}{p!q!}\mathbb{I}_{p}\bigg(\int_{[0,1]^{q}}\phi_{k}(t_{1},\cdots,t_{k})dt_{p+1}\cdots dt_{k}\bigg)$
@@
Assume (A2). Then we have $\mathrm{Tr}^{j}{\psi}_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $j=0,1,...,\lfloor\frac{m}{2}\rfloor$. Moreover,
@@
$\|\mathrm{Tr}^{j}{\phi}_{m}$ $\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\phi_{m}(s_{1},\cdots,s_{2j},t_{1},\cdots,t_{m-2j})$
$\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m}(s_{1}^{\prime},\cdots,s_{2j}^{\prime},t_{1}^{\prime},\cdots,t_{m-2j}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
First, we split $0,1]^{2m}$ into $(2m)!$ regions such that in each region, the numbers $\{t_{i},t_{i}^{\prime},s_{j},s_{j}^{\prime}:i=1,\dots,m-2j,j=1,\dots,2j\}$ are listed in an increasing order. To get an upper bound, we replace, for instance, each $|t_{i}-t_{i}^{\prime}|^{2H-2}$ by $|t_{i}^{*}-t_{i}^{\star}|^{2H-2}$ where $t^{*}_{i}:=\max\{t_{i},t_{i}^{\prime}\}$ and $t_{i}^{\star}$ is the largest number that is smaller than $t^{*}_{i}$ (it could be some $t_{j}$, $t^{\prime}_{j}$, $s_{j}$ or $s^{\prime}_{j}$), and we perform the same operation for all the other $|s_{2i}-s_{2i-1}|^{2H-2}$, $|s^{\prime}_{2i}-s^{\prime}_{2i-1}|^{2H-2}$ and the factors in $\phi_{m}$. Then on each region, the above integral is smaller than an integral with $2m$ variables in Lemma A.5 with $m$ exponents equal to ${\alpha}-1$ and $m$ exponents equal to $2H+{\alpha}-3$. Noting that $2H+\alpha-3>-1$ under (A2), we get by Lemma A.5
@@
When $k>0$, by denoting $\boldsymbol{s}:=(s_{1},\cdots,s_{2j})$, $\boldsymbol{t}:=(t_{1},\cdots,t_{m-2j})$, $\boldsymbol{x}:=(x_{1},\cdots,x_{k})$ and $\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime}$ correspondingly, we have
$\quad\|\mathrm{Tr}^{j}{\psi}_{m,k}\|^{2}_{|\mathcal{H}|^{\otimes(m-2j)}}=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\psi_{m,k}(\boldsymbol{s},\boldsymbol{t})$
$\qquad\qquad\qquad\qquad\qquad\times\int_{[0,1]^{2j}}\psi_{m,k}(\boldsymbol{s}^{\prime},t_{1}^{\prime},\boldsymbol{t}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
$=\int_{[0,1]^{2(m-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\bigg(\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(\boldsymbol{s},\boldsymbol{t},\boldsymbol{x})\mathrm{d}\boldsymbol{x}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{[0,1]^{2j}}\int_{[0,1]^{k}}\phi_{m+k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime})\mathrm{d}\boldsymbol{x^{\prime}}\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)$
$\leq\int_{[0,1]^{2(m+k-2j)}}\prod_{i=1}^{m-2j}|t_{i}-t_{i}^{\prime}|^{2H-2}\mathrm{d}t_{i}\mathrm{d}t_{i}^{\prime}\prod_{p=1}^{k}|x_{p}-x_{p}^{\prime}|^{2H-2}\mathrm{d}x_{p}\mathrm{d}x_{p}^{\prime}\bigg(\int_{[0,1]^{2j}}\phi_{m+k}(\boldsymbol{s},\boldsymbol{t},\boldsymbol{x})$
$\quad\times\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}\mathrm{d}s_{2i}}{|s_{2i}-s_{2i-1}|^{2-2H}}\int_{[0,1]^{2j}}\phi_{m+k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime},\boldsymbol{x}^{\prime})\prod\limits_{i=1}^{j}\frac{\mathrm{d}s_{2i-1}^{\prime}\mathrm{d}s_{2i}^{\prime}}{|s_{2i}^{\prime}-s_{2i-1}^{\prime}|^{2-2H}}\bigg)=\|\mathrm{Tr}^{j}{\phi}_{m+k}\|^{2}_{|\mathcal{H}|^{\otimes(m+k-2j)}}.$
@@
Assuming (A2), Lemma 4.1 yields $\interleave\psi_{m,k}\interleave_{m}<\infty$} (see (2.11)) and hence $\mathbb{I}_{m}(\psi_{m,k})$ is well defined due to Hu–Meyer formula (2.10). On the other hand, $\alpha+2H>2$ is the optimal condition for $\interleave\psi_{m,k}\interleave_{m}<\infty$}. Indeed, for $k=0,m=2$ and $j=1$, we have
$\mathrm{Tr}^{1}\phi_{2}=2\int_{0<t_{1}<t_{2}<1}\frac{C_{\alpha}^{2}}{t_{1}^{1-{\alpha}}(t_{2}-t_{1})^{1-{\alpha}}}(t_{2}-t_{1})^{2H-2}dt_{1}dt_{2},$
@@
Recalling $Z_{\hat{\beta},\hat{h}}$ from (4.1), we have
$\|Z_{\hat{\beta},\hat{h}}\|_{2}$ $\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\|\mathbb{I}_{m}(\psi_{m,k})\|_{2}\leq\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{\hat{\beta}^{m}|\hat{h}|^{k}}{m!k!}\sqrt{2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}}$
@@
where the second inequality follows from Lemma 4.4 below and the last inequality is due to (4.2) and (4.3) above.
@@
$\mathbb{E}[\mathbb{I}_{m}(\phi_{m})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\phi}_{2m}\quad\text{and}\quad\mathbb{E}[\mathbb{I}_{m}(\psi_{m,k})^{2}]\leq 2^{m}(m+1)!\mathrm{Tr}^{m}{\psi}_{2m,2k}.$
@@
For the first inequality, Hu–Meyer formula (2.10) implies
$\mathbb{E}[|\mathbb{I}_{m}(\phi_{m})|^{2}]$ $=\sum_{k=0}^{[\frac{m}{2}]}\frac{1}{(m-2k)!}\left(\frac{m!}{k!2^{k}}\right)^{2}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}\binom{m}{2k}\binom{2k}{k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.4)
$\leq\sum_{k=0}^{[\frac{m}{2}]}\frac{m!}{2^{2k}}2^{m+2k}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=2^{m}m!\sum_{k=0}^{[\frac{m}{2}]}\|\mathrm{Tr}^{k}\phi_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}},$